# A metrological framework for uncertainty evaluation in machine learning classification models

### v2

@@ 3.2 Expressions of uncertainty
The first point to note is the distinction made in the literature e.g., [mari_foundations_2020, shirmohammadi_measurement_2024], between the confidence in the prediction of an ML classification model, and the predictive uncertainty. Given a probability vector $\mathbf{p}=[p_{1},,\dots,p_{K}]$ of an ML classification model $y=f(x),\,y\in\mathcal{C}_{K}$, the classification of the output is typically given by the mode of the PMF:
@@
The confidence in the prediction can then be expressed as $p_{\mathrm{mode}}$, which is the probability that $y$ is the modal class $c_{\mathrm{mode}}$, also known as the modal probability. Note that this is distinct from a confidence interval which is used to express uncertainty for quantitative variables in the GUM and the VIM.
Using the modal probability as a form of expression has advantages in the sense that it is directly associated with the model prediction, and as such aligns with the language for expressing uncertainty of quantitative variables [JCGMGUM]. For example, the output of a classification model $\widehat{y}=f(x)$ can be written as the pair $(c_{\mathrm{mode}},p_{\mathrm{mode}})$, where $p_{\mathrm{mode}}$ is the confidence in the prediction that $\widehat{y}=c_{\mathrm{mode}}$. This language can be extended [mari_foundations_2020] to include multiple classes, i.e., the ML classification model predicts $\widehat{y}=\{c_{1},c_{2}\}$ with probability $p=p_{1}+p_{2}$. However, confidence is not the same as uncertainty. A confidence equal to $1$ would express complete certainty in the output of a classification model, thus the uncertainty should be 00.
@@ 3.2.1 Properties of an uncertainty statistic
In general, a statistic $F$ expressing the uncertainty associated with a nominal variable $y$ with PMF $\mathbf{p}$, should satisfy the following properties (adapted from [wilcox_indices_1973, kvalseth_variation_2011]):
@@
    $F$ gives a positive real scalar value: $F(\mathbf{p})\in[0,\infty)$.
@@
    $F$ gives 00 when there is no uncertainty. That is, $F(\mathbf{p}^{1})=0$, where
    $p^{1}_{k}=\begin{cases}1,&k=\mathrm{mode},\\ 0,&\text{otherwise.}\end{cases}$
@@
    $F$ is maximum when there is maximal uncertainty. This is when all probabilities are equal:
    $F(\mathbf{p})\leq F(\mathbf{p}_{\text{unif}}),$
@@
In summary, a (normalised) statistic $F$ expressing the uncertainty associated with the output probabilities of an ML classification model is minimum (zero) when the model predicts equal probabilities to all classes, and is maximum (one) when the model predicts a single class with $100\,\%$ probability.
In the following analysis, we will consider only normalised statistics that satisfy the above properties, for a fair comparison between expressions of uncertainty. As such, we do not consider statistics that quantify different aspects of uncertainty, such as the expected difference of information [bogaert_information-based_2017], which has been gaining popularity in land cover classification (see case study in Section 4.1). This statistic requires identification of a reference class, which fails on property 3, since the reference probability is minimum for a uniform distribution.
3.2.2 Information entropy
The information (or Shannon) entropy (not to be confused with thermodynamic entropy) is an expression of information content, originally motivated to describe the theory of signal communication [shannon_mathematical_1948].
@@ 3.2.2 Information entropy
Information entropy is commonly used in the existing literature as a statistic for expressing the uncertainty in nominal variables [mari_toward_2017, wilcox_indices_1973, possolo_statistical_2014], as well as the recent international standard [ISO33406]. It is also the most commonly used statistic in the ML community for summarising uncertainty in predictions of classification models [kendall2017uncertainties, shirmohammadi_measurement_2024, hullermeier2021aleatoric]. The principle of maximum entropy (uncertainty) is also used in Bayesian inference (see Section 4.1.3) for selection of prior probabilities [jaynes_prior_1968].
Mari [mari_foundations_2020] suggested an extension to information entropy given (in normalised form) by
@@
Such a transformation can be performed to any expression of uncertainty $F(\mathbf{p})$ for nominal properties to produce a new statistic $F^{*}(\mathbf{p})=(K^{F(\mathbf{p})}-1)/(K-1)$, which also satisfies the properties in Section 3.2.1. This transformation has the property $F^{*}(\mathbf{p})\leq F(\mathbf{p})\,\forall\,\mathbf{p}$, and so will lower the variation in the values of the new statistic (see Figures 5 and 7). This allows the statistic to be more robust against small perturbations of the PMF.
One alternative variant on the information entropy is called the $\alpha$-quadratic entropy [pal_measuring_1994], defined as
@@
If $\alpha$ is close to zero, this statistic is not very sensitive to changes in $\mathbf{p}$, but for $\alpha$ close to one, this statistic has higher sensitivity [low_impact_2013], and has been used in evaluating uncertainties in land cover classification models [low_impact_2013, Bogaert2017]. We consider the case when $\alpha=1$ in Section 3.2.6.
3.2.3 Wilcox’s variation ratio
Wilcox’s variation ratio (WVR) [wilcox_indices_1973], also known as the mean deviation from the mode, is a natural extension of the concept of confidence in a classification model prediction to an expression of uncertainty, defined by
$u_{\mathrm{WVR}}(\mathbf{p})=1-\dfrac{Kp_{\mathrm{mode}}-1}{K-1}.$ (3.12)
The WVR expresses the remaining uncertainty after taking into account the modal probability $p_{\mathrm{mode}}$ of the classification result. However, it does not take into account the entire PMF, so is ignorant of variation in probabilities of the remaining classes. This issue is only relevant in the case of non-binary classification problems, such as that described in Section 4.1.
3.2.4 Universal variation ratio
The WVR is undefined in the case of multimodal distributions, where more than one class is equally likely. The universal variation ratio (UVR) [rubia_measures_2024] is an extension of the WVR that takes account of multimodal distributions, defined as
$u_{\mathrm{UVR}}(\mathbf{p})=\dfrac{K^{2}}{K^{2}-1}\left(1-\dfrac{p_{\mathrm{mode}}}{c}\right),$ (3.13)
where $c$ is the number of modes. It can be shown [rubia_measures_2024] that $u_{\mathrm{UVR}}(\mathbf{p})\leq u_{\mathrm{WVR}}(\mathbf{p})$ for all unimodal distributions.
3.2.5 Standard deviation from the mode
Both the WVR and UVR take account of the modal probability $p_{\mathrm{mode}}$ only, but not the distribution among the remaining classes, which might be important to consider. Although information entropy takes account of all class probabilities, there is no notion of “distance” from the central tendency. Motivated by finding an equivalent to the standard deviation about the mean for quantitative variables, a standard deviation from the mode (SDM) for nominal variables [kvalseth_measuring_1988] can be defined as
$u_{\mathrm{SDM}}(\mathbf{p})=1-\sqrt{\dfrac{\sum_{k=1}^{K}(p_{\mathrm{mode}}-p_{k})^{2}}{K-1}}.$ (3.14)
One advantage of using the SDM is that for a large number $n$ of samples used to estimate the output PMF, such as the output trees of random forests, the SDM is normally distributed [kvalseth_measuring_1988] $\widehat{u}_{\mathrm{SDM}}\sim\mathcal{N}(u_{\mathrm{SDM}}(\mathbf{p}),\sigma_{\mathrm{SDM}}^{2}(\mathbf{p}))$ with
$\sigma_{\mathrm{SDM}}^{2}(\mathbf{p})=\dfrac{p_{\mathrm{mode}}(1-Kp_{\mathrm{mode}})^{2}+\sum_{k=1}^{K}p_{k}(p_{\mathrm{mode}}-p_{k})^{2}}{n(K-1)^{2}(1-u_{\mathrm{SDM}}(\mathbf{p}))^{2}}-\dfrac{(1-u_{\mathrm{SDM}}(\mathbf{p}))^{2}}{n}.$ (3.15)
This result allows the use of confidence intervals for the statistic.
3.2.6 Index of qualitative variation
The index of qualitative variation (IQV) [wilcox_indices_1973], also known as the quadratic score [glasziou_test_1989], is another common expression of uncertainty that is widely accepted in the social sciences [rubia_measures_2024]. It is defined in normalised form as
$u_{\mathrm{IQV}}(\mathbf{p})=\dfrac{K}{K-1}\left(1-\sum_{k=1}^{K}p_{k}^{2}\right).$ (3.16)
@@ 3.2.6 Index of qualitative variation
3.2.7 Coefficient of nominal variation
A variant of the IQV proposed in [kvalseth_coefficients_1995] is the coefficient of nominal variation (CNV) defined as
@@ 3.2.7 Coefficient of nominal variation
This transformation ensures that $u_{\mathrm{CNV}}(\mathbf{p})\leq u_{\mathrm{IQV}}(\mathbf{p})$ for all distributions, with equality for $\mathbf{p}^{1}$ and $\mathbf{p}_{\text{unif}}$. In terms of the normed distance $d(\mathbf{p},\mathbf{q})=||\mathbf{p}-\mathbf{q}||_{2}$ between two PDFs $\mathbf{p}$ and $\mathbf{q}$, the CNV can be rewritten as
$u_{\mathrm{CNV}}(\mathbf{p})=1-\dfrac{d(\mathbf{p},\mathbf{p}_{\text{unif}})}{d(\mathbf{p}^{1},\mathbf{p}_{\text{unif}})},$
giving the CNV a natural interpretation in terms of the distance of the PMF from a uniform distribution.
@@ 3.2.8 Comparing statistics of uncertainty
The statistics described in Section 3.2 have various advantages and disadvantages. In this section we compare their properties and performance on a few example predictive probability distributions. One particular example in binary classification is the PMF $\mathbf{p}_{A}$ shown in Figure 3. Since $\mathbf{p}_{A}$ is halfway between total certainty and maximal uncertainty, it was suggested in [kvalseth_variation_2011] that the statistic of uncertainty should also be halfway between the extreme values, i.e., 0.5. We also consider the multi-class PMFs $\mathbf{p}_{B}$ and $\mathbf{p}_{C}$, which have the same model prediction $\widehat{y}$, with the same confidence (modal probability) of 0.5 (see Figure 3).
@@
Table 2 compares the various expressions of uncertainty for nominal properties. The first point to note is the result that $F(\mathbf{p}_{B})\geq F(\mathbf{p}_{C})$ for all $F$ considered, even though it seems more obvious which is the best prediction of the classification model in $\mathbf{p}_{B}$. That is because the statistics are expressing the certainty that the mode can either be one of two possibilities (i.e. $(p_{C})_{1}+(p_{C})_{2}=0.96$ for $\mathbf{p}_{C}$), whereas there is more spread in the distribution of $\mathbf{p}_{B}$.
If we expect the uncertainty of $\mathbf{p}_{A}$ to be 0.5, then we see that WVR, SDM and CNV satisfy this condition. If information about the modal probability is to be used explicitly and information about the entire distribution is to be used, then SDM is the only option. However, the SDM shows little difference in uncertainty between $\mathbf{p}_{B}$ and $\mathbf{p}_{C}$, where a larger difference from studying Figure 3 might be expected to be seen. In this case, CNV might be the more appropriate choice, but at the cost of not explicitly using the modal probability.
We specify in Table 2 whether a particular statistic is well-defined (Y) or not well-defined (N) in the pathological multimodal case where two or more classes share precisely the same modal probability. However, the definition of these statistics can be trivially extended to the multimodal case by arbitrarily choosing the mode to be any of the maximum probability classes, this choice not affecting the value of the statistic.
@@
Statistic Uses $p_{\mathrm{mode}}$ Uses full PMF Multimodal $\mathbf{p}_{A}$ $\mathbf{p}_{B}$ $\mathbf{p}_{C}$
$H$ (3.9) N Y Y 0.81 0.84 0.50
$H^{*}$ (3.10) N Y Y 0.75 0.69 0.29
WVR (3.12) Y N N 0.50 0.60 0.60
UVR (3.13) Y N Y 0.33 0.51 0.51
SDM (3.14) Y Y N 0.50 0.60 0.56
IQV (3.16) N Y Y 0.75 0.84 0.65
CNV (3.17) N Y Y 0.50 0.60 0.40
@@ 3.3 Uncertainty propagation with nominal properties
In Section 3.2, we considered various ways in which the uncertainty in nominal properties can be expressed through a summary statistic. However, to propagate such uncertainty through a chain of measurement models, it is more appropriate to consider the full PMF. For quantitative properties, it is more difficult to propagate the full PDF analytically due to intractable integrals, and it is commonly the case that the full PDF is unknown and modelled using simpler forms such as assuming a Gaussian distribution. In such cases, summary statistics, such as the standard deviation, are propagated using the law of propagation of uncertainty [JCGMGUM]. Alternatively, where the PDF is well known, sampling techniques such as Monte Carlo (MC) can propagate the full PDF [JCGM101].
In the case of a PDF having an analytical form, such as a Gaussian or rectangular distribution, the full PDF can be directly inferred from summary statistics, such as the standard deviation. On the other hand, the summary statistics for nominal properties described in Section 3.2 cannot be used to infer the PMF except in a few very niche examples. This property also supports the use of the full PMF in uncertainty propagation in traceability chains that involve nominal properties.
@@
[…]
@@
We can also compute the variance of the output $z$:
@@
$=\sum_{k=1}^{K}p_{k}(\text{Var}[g_{k}(\mathbf{x})]+\mu_{k}^{2})-\sum_{j,k=1}^{K}p_{j}p_{k}\mu_{j}\mu_{k}$
$=\sum_{k=1}^{K}p_{k}\sigma_{k}^{2}+\sum_{j,k=1}^{K}p_{j}p_{k}(\mu_{k}^{2}-\mu_{j}\mu_{k}),$
where $\sigma_{k}^{2}=\text{Var}_{\mathbf{x}\sim q_{k}(\mathbf{x})}[g_{k}(\mathbf{x})]$ can be evaluated using the methods described in [JCGMGUM] for computing expressions of uncertainty in quantitative variables.
For this simple example, the moments of the measurand given nominal property inputs can be determined analytically. More generally, MC techniques can be applied by sampling directly from the PMF to determine the output PDF $p(z)$. Proceeding in this way is relevant to measurement applications that depend upon a prior classification, such as in determining particle concentration from counting measurements [sipkens_tutorial_2023].

### v3

@@ 3.2.1 Uncertainty and confidence
A distinction is made in the literature e.g., [35, 40], between the confidence in the prediction of an ML classification model, and the predictive uncertainty. Given a probability vector $\mathbf{p}=[p_{1},,\dots,p_{K}]$ of an ML classification model $y=f(x),\,y\in\mathcal{C}_{K}$, the classification of the output is typically given by the mode of the PMF, defined as the modal class $\widehat{c}$. The confidence in the prediction can then be expressed as modal probability $\widehat{p}$, which is the probability that $y$ is the modal class. Note that this notion of confidence is distinct from a confidence interval which is one way of expressing uncertainty for quantitative variables in the GUM and the VIM.
@@ 3.2.2 Properties of an uncertainty statistic
In general, a statistic $F$ expressing the uncertainty associated with a nominal variable $y$ with PMF $\mathbf{p}$, should satisfy the following properties (adapted from [41, 39]):
@@
    $F$ gives a non-negative real scalar value: $F(\mathbf{p})\in[0,\infty)$.
@@
    $F$ gives 0 if and only if there is no uncertainty. That is, $F(\mathbf{p}^{1})=0$ if and only if the elements of $\mathbf{p}^{1}$ take the form
    $p^{1}_{k}:=\begin{cases}1&c_{k}=\widehat{c},\\ 0&\text{otherwise,}\end{cases}\quad k=1,\dots,K.$ (3.9)
@@
    $F$ is maximum when there is maximal uncertainty. This is when all probabilities are equal, or equivalently when all classes could be the mode:
    $F(\mathbf{p})\leq F(\mathbf{p}_{\text{unif}}),$ (3.10)
@@
In summary, a (normalised) statistic $F$ expressing the uncertainty associated with the output probabilities of an ML classification model is minimum (zero) when the model predicts a single class with $100\,\%$ probability, and is maximum (one) when the model predicts equal probabilities to all classes.
In the following analysis, we will consider only normalised statistics that satisfy the above properties, for a fair comparison between expressions of uncertainty (see Section 3.2.4). As such, we do not consider statistics that quantify different aspects of uncertainty, such as the expected difference of information to a reference class [42], which has been gaining popularity in land cover classification (see case study in Section 4.1). This statistic fails on property 3, since the statistic is minimum for a uniform distribution, whereas it should be maximum if it is to represent uncertainty.
3.2.3 Statistics of dispersion
In this section, we give an overview of some of the statistics of dispersion that have been proposed.
Wilcox’s variation ratio
Wilcox’s variation ratio (WVR) [41], also known as the mean deviation from the mode, is a natural extension of the concept of confidence (see Section 3.2.1) in a classification model prediction to an expression of uncertainty, defined by
$u_{\mathrm{WVR}}(\mathbf{p})=1-\dfrac{K\widehat{p}-1}{K-1},$ (3.12)
For unimodal distributions only, the WVR can be expressed in terms of the expected distance from the mode. Thus the WVR is analogous to the standard deviation for continuous variables. See Appendix A for a derivation of WVR in terms of expected distance from the mode.
For binary classification ($K=2$), expression (3.12) reduces to
[…]
The WVR expresses the remaining uncertainty after taking into account the modal probability $\widehat{p}$ of the classification result. However, it does not take into account the entire PMF, and and in particular variation in probabilities of the remaining classes. This issue is only relevant in the case of non-binary classification problems, such as the one described in Section 4.1.
Universal variation ratio
The universal variation ratio (UVR) [43] is an extension of the WVR that explicitly takes account of multimodal distributions, defined as
$u_{\mathrm{UVR}}(\mathbf{p})=\dfrac{K^{2}}{K^{2}-1}\left(1-\dfrac{\widehat{p}}{m}\right),$ (3.14)
where $m$ is the number of modes. This extension allows expression of the UVR in terms of the expected distance from the mode in the multimodal case (see Appendix A for a derivation), which cannot be achieved for the WVR.
By explicitly including the number of modes, the UVR is discontinuous as $\widehat{p}$ varies. This becomes clear in binary classification, where $m=1$ in all cases except for a uniform distribution when $m=2$. Therefore, in binary classification expression (3.14) reduces to
$u_{\mathrm{UVR}}(\mathbf{p})=\begin{cases}\frac{4}{3}(1-\widehat{p})&\widehat{p}>0.5,\\ 1&\widehat{p}=0.5.\end{cases}$ (3.15)
It can be shown [43] that $u_{\mathrm{UVR}}(\mathbf{p})\leq u_{\mathrm{WVR}}(\mathbf{p})$ for all unimodal distributions.
Standard deviation from the mode
Both the WVR and UVR take account of the modal probability $\widehat{p}$ only, but not the distribution among the remaining classes, which might be important to consider. Motivated by finding a statistic based on the deviation from the mode, which takes account of all values in the PMF, Kvålseth [44] defined the standard deviation from the mode (SDM) for nominal variables. In normalized form this standard deviation is
$u_{\mathrm{SDM}}(\mathbf{p})=1-\sqrt{\dfrac{\sum_{k=1}^{K}(\widehat{p}-p_{k})^{2}}{K-1}}.$ (3.16)
For binary classification, expression (3.16) reduces to (see Appendix B):
[…]
and so in the binary case WVR and SDM are equivalent.
One advantage of using the SDM is that for a large number $n$ of samples used to estimate the output PMF, such as the output trees of random forests, the SDM is normally distributed in the unimodal case [44] $\widehat{u}_{\mathrm{SDM}}\sim\mathcal{N}(u_{\mathrm{SDM}}(\mathbf{p}),\sigma_{\mathrm{SDM}}^{2}(\mathbf{p}))$ with
$\sigma_{\mathrm{SDM}}^{2}(\mathbf{p})=\dfrac{\widehat{p}(1-K\widehat{p})^{2}+\sum_{k=1}^{K}p_{k}(\widehat{p}-p_{k})^{2}}{n(K-1)^{2}(1-u_{\mathrm{SDM}}(\mathbf{p}))^{2}}-\dfrac{(1-u_{\mathrm{SDM}}(\mathbf{p}))^{2}}{n}.$ (3.18)
This result allows the use of confidence intervals for the statistic.
Entropy
The information (or Shannon) entropy (not to be confused with thermodynamic entropy) is an expression of information content, originally motivated to describe the theory of signal communication [45].
@@ 3.2.3 Statistics of dispersion
In the binary classification case, expression (3.19) reduces to the original Shannon entropy in bits
[…]
Information entropy is commonly used in the existing literature as a statistic for expressing the uncertainty in nominal variables [46, 41, 47], as well as the recent international standard [32]. It is also the most commonly used statistic in the ML community for summarising uncertainty in predictions of classification models [19, 40, 15]. The principle of maximum entropy (uncertainty) is also used in Bayesian inference (see Section 4.1.3) for selection of prior probabilities [48].
Mari [35] suggested an extension to information entropy given (in normalised form) by
@@
Such a transformation can be performed to any expression of uncertainty $F(\mathbf{p})$ for nominal properties to produce a new statistic $F^{*}(\mathbf{p})=(K^{F(\mathbf{p})}-1)/(K-1)$, which also satisfies the properties in Section 3.2.2. This transformation has the property $F^{*}(\mathbf{p})\leq F(\mathbf{p})\,\forall\,\mathbf{p}$, and so will lower the variation in the values of the new statistic (see Tables 45 and 6). This allows the statistic to be more robust against small perturbations of the PMF.
One alternative variant on the information entropy is called the $\alpha$-quadratic entropy [49], defined as
@@
If $\alpha$ is close to zero, this statistic is not very sensitive to changes in $\mathbf{p}$, but for $\alpha$ close to one, this statistic has higher sensitivity [50], and has been used in evaluating uncertainties in land cover classification models [50, 42].
Both of the statistics given in (3.21) and (3.22) reduce to statistics of dispersion about the mode in the binary case, but we omit the details for the sake of brevity.
Index of qualitative variation
The index of qualitative variation (IQV) [41], also known as the quadratic score [51], is another common expression of uncertainty that is widely accepted in the social sciences [43]. It is defined in normalised form as
$u_{\mathrm{IQV}}(\mathbf{p})=\dfrac{K}{K-1}\left(1-\sum_{k=1}^{K}p_{k}^{2}\right),$ (3.23)
which for binary classification reduces to
[…]
@@
Coefficient of nominal variation
A variant of the IQV proposed in [52] is the coefficient of nominal variation (CNV) defined as
@@
[…]
$u_{\mathrm{CNV}}(\mathbf{p})=1-\dfrac{d(\mathbf{p},\mathbf{p}_{\text{unif}})}{d(\mathbf{p}^{1},\mathbf{p}_{\text{unif}})},$ (3.26)
giving the CNV a natural interpretation in terms of the distance of the PMF from a uniform distribution. For binary classification, expression (3.25) reduces to (see Appendix B):
[…]
which is equivalent to the WVR and the SDM.
@@ 3.2.4 Comparing statistics of uncertainty
The statistics described in Section 3.2 have various advantages and disadvantages. In this section we compare their properties and performance on a few example predictive probability distributions. One particular example in binary classification is the PMF $\mathbf{p}_{A}$ shown in Figure 3. Since $\mathbf{p}_{A}$ is halfway between total certainty and maximal uncertainty, i.e. $\widehat{p}_{A}=0.75$, it was suggested in [39] that the statistic of uncertainty should also be halfway between the extreme values, i.e., 0.5. We also consider the multi-class PMFs $\mathbf{p}_{B}$ and $\mathbf{p}_{C}$, which have the same model prediction $\widehat{y}$, with the same confidence (modal probability) of 0.5. The example PMFs are visualised in Figure 3 for comparison.
@@
Table 2 compares the various expressions of uncertainty for nominal properties. The first point to note is the result that $F(\mathbf{p}_{B})\geq F(\mathbf{p}_{C})$ for all $F$ considered, even though at first glance it might appear more obvious which is the best prediction of the classification model in $\mathbf{p}_{B}$. This result is due to the fact that the statistics are expressing the certainty that the mode can either be one of two possibilities in $\mathbf{p}_{C}$ (i.e. $(p_{C})_{1}+(p_{C})_{2}=0.96$), as the distribution is peaked about these two possible classes. Whereas for $\mathbf{p}_{B}$, the probability distribution is spread over all classes, leading to more uncertainty in the output class.
If we expect the uncertainty of $\mathbf{p}_{A}$ to be 0.5, then we see that WVR, SDM and CNV satisfy this condition. In fact, WVR, SDM and CNV have the same expressions for binary classification (see Appendix B). These three statistics also have the same values for $\mathbf{p}_{B}$ because expressions (3.16) and (3.25) simplify to (3.12) when all the non-modal classes have the same probability (see Appendix B). Thus, we only see a difference in values for $\mathbf{p}_{C}$, with WVR and CNV giving the highest and lowest values respectively.
If we wish the statistic to use explicitly the modal probability (or confidence), as well as information about the entire distribution, then SDM is the only option. However, the SDM shows little difference in uncertainty between $\mathbf{p}_{B}$ and $\mathbf{p}_{C}$, where a larger difference from studying Figure 3 might be expected. In this case, CNV might be the more appropriate choice, but at the cost of no longer explicitly using information about the mode.
@@
Statistic Uses confidence Uses full PMF $\mathbf{p}_{A}$ $\mathbf{p}_{B}$ $\mathbf{p}_{C}$
WVR (3.12) Y N 0.50 0.60 0.60
UVR (3.14) Y N 0.33 0.51 0.51
SDM (3.16) Y Y 0.50 0.60 0.56
$H$ (3.19) N Y 0.81 0.84 0.50
$H^{*}$ (3.21) N Y 0.75 0.69 0.29
IQV (3.23) N Y 0.75 0.84 0.65
CNV (3.25) N Y 0.50 0.60 0.40
@@ 3.3 Uncertainty propagation with nominal properties
In Section 3.2, we considered various ways in which the uncertainty in nominal properties can be expressed through a summary statistic. However, to propagate such uncertainty through a chain of measurement models, it is more appropriate to consider the full PMF. For quantitative properties, it is more difficult to propagate the full PDF analytically due to intractable integrals, and it is commonly the case that the full PDF is unknown and modelled using simpler forms such as assuming a Gaussian distribution. In such cases, summary statistics, such as the standard deviation, are propagated using the law of propagation of uncertainty [21]. Alternatively, where the PDF is well known, sampling techniques such as Monte Carlo (MC) can propagate the full PDF [22].
In the case of a PDF having an analytical form, such as a Gaussian or rectangular distribution, the full PDF can be directly inferred from summary statistics, such as the standard deviation. On the other hand, the summary statistics for nominal properties described in Section 3.2 cannot be used to infer the PMF except in simplistic examples such as binary classification. It is therefore recommended to use the full PMF in uncertainty propagation through multistage measurement models that involve nominal properties.
@@
$\mathbb{E}[z]=\sum_{k=1}^{K}p_{k}\int_{\mathcal{X}_{k}}\mathrm{d}\mathbf{x}\,q_{k}(\mathbf{x})g(\mathbf{x},c_{k})=\sum_{k=1}^{K}p_{k}\int_{\mathcal{X}_{k}}\mathrm{d}\mathbf{x}\,q_{k}(\mathbf{x})g_{k}(\mathbf{x})=\sum_{k=1}^{K}p_{k}\mu_{k},$ (3.29)
@@
We can also compute the variance of the output $z$ in terms of summary statistics of the functions $g_{k}(\mathbf{x})$:
@@
$=\sum_{k=1}^{K}p_{k}(\sigma_{k}^{2}+\mu_{k}^{2})-\sum_{j,k=1}^{K}p_{j}p_{k}\mu_{j}\mu_{k}$ (3.30)
where $\sigma_{k}^{2}=\text{Var}_{\mathbf{x}\sim q_{k}(\mathbf{x})}[g_{k}(\mathbf{x})]$ can be evaluated using the methods described in [21] for computing expressions of uncertainty in quantitative variables.
For this simple example, the moments of the measurand given nominal property inputs can be determined analytically. More generally, MC techniques can be applied by sampling directly from the PMF to determine the output PDF $p(z)$. Proceeding in this way is relevant to measurement applications that depend upon a prior classification, such as in determining particle concentration from counting measurements [53].